R=(7+4*root3)^{2n}=I+f
find R(1-f)
ย 

Grade:8

1 Answers

Arun Kumar IIT Delhi
8 years ago
Hello Student,
G=(7-4*root3)^{2n}=Even integer.
I+f+G=an even integer.
f+G=1=>G=1-f
=>R(1-f)=RG=1
Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty